OpenStudy (anonymous):

use the quadratic formula to solve x^2+3x-5=0

OpenStudy (anonymous):

the constant in front of the x^2 is a the constant in front of x is b and the constant is c from here -b(+/-)sqrt(b^2-4ac)/2a

OpenStudy (anonymous):

so you have two solutions (-3+sqrt(3^2-4*1*5))/(2*1)

OpenStudy (anonymous):

and (-3-sqrt(3^2-4*1*5))/(2*1)

OpenStudy (anonymous):

The number in the sqrt is negative so you have a non-real answer

OpenStudy (anonymous):

@hera474 has the right answer, but the equation is a little hard to read so I'll rewrite it using the equation tool:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

@tcarroll010 Thanks i didn't know I could do that :)

OpenStudy (anonymous):

thank you all bout i solved and got -1.925 is that right?

OpenStudy (anonymous):

Have you learned about imaginary numbers yet?

OpenStudy (anonymous):

The equation tool is handy, useful, and fun. It uses LatX document preparation language and it sets that up all behind the scenes. I suggest to anyone to experiment with it.

OpenStudy (anonymous):

It's either no solution or you have to give it in complex for.

OpenStudy (anonymous):

And this problem does not have imaginary numbers for the solution because "a" and "c" have opposite signs. Whenever you have that, you get real solutions.

OpenStudy (anonymous):

Wow... That was stupid of me.

OpenStudy (anonymous):

The number in the radical will be positive. It was not stupid of you, it was just an oversight. Everyone does that from time to time. Keep up with your explanation, you're doing fine. You just went slightly off for a second.

OpenStudy (anonymous):

can someone help me solve this ?

OpenStudy (anonymous):

a=1, b=3, and c=-5, so start by substituting those values into the formula you were given by us. The a, b, and c come from: ax^2 + bx + c = 0 so you might want to first match those a, b, and c up from your equation to see where we got those from.

OpenStudy (anonymous):

So, @bri4life14 , you have been given some good help here. Where do you stand with your work on this? Show us some of your work now.

OpenStudy (anonymous):

Dang this equation thing takes a while

OpenStudy (anonymous):

im stuck :( i put everything into the formula but i dont know where to start

OpenStudy (anonymous):

\[(-3-\sqrt{3^2-4*1*-5})/(2*1)\] and \[(-3+\sqrt{3^2-4*1*-5})/(2*1)\]

OpenStudy (anonymous):

\[(-3+\sqrt(9+20))/2\] and \[(-3-\sqrt(9+20))/2\]

OpenStudy (anonymous):

You can get the answer if you plug it into a calculator.

OpenStudy (anonymous):

Do you understand it now?

OpenStudy (anonymous):

yes im trying to do this with you ....how did u get 9

OpenStudy (anonymous):

3 squared is 9

OpenStudy (anonymous):

oh sorry and i got -5.692

OpenStudy (anonymous):

you should get two answers one from each equation.